OpenStudy (anonymous):

Male Female Total Income over $55,000 475 375 850 Income below $55,000 75 75 150 Total 550 450 1,000 Based on these data, are "being male" and "earning over $55,000" independent events? No, P(being male | the person earns over $55,000) ≠ P(being male) No, P(being male | the person earns over $55,000) = P(being male) Yes, P(being male | the person earns over $55,000) = P(being male) Yes, P(being male | the person earns over $55,000) ≠ P(being male)

OpenStudy (anonymous):

@amistre64 @rational

OpenStudy (anonymous):

is it B?

OpenStudy (anonymous):

A. No, P(being male...

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

@_Smile_

jimthompson5910 (jim_thompson5910):

compute the probability P(being male | the person earns over $55,000) and the probability P(male)

jimthompson5910 (jim_thompson5910):

P(being male | the person earns over $55,000) = x/y x = number of males who earn over $55,000 y = number of people who earn over $55,000

jimthompson5910 (jim_thompson5910):

P(male) = z/w z = number of males w = number of people total

jimthompson5910 (jim_thompson5910):

Rule: Let A and B be two events. They are said to be independent if and only if P(A|B) = P(A) or P(B|A) = P(B) otherwise, they are dependent

OpenStudy (amistre64):

if the probability of being male is the same regardless of the event that takes place, then the probability is independant of any given event.

OpenStudy (anonymous):

I tagged you in something DoShKa_SyRiA

OpenStudy (amistre64):

if the probability of male, given the event over$55 is the same as the probability of male, given the totality of events then the probability of male does not depend on a specific event.

OpenStudy (anonymous):

so 550/850 ?

OpenStudy (amistre64):

475 out of 850 is not the same as 550 out of 1000 so we know they are not

OpenStudy (amistre64):

Male Female Total Income over $55,000 475 375 850 Income below $55,000 75 75 150 Total 550 450 1,000 Male Female Total Income over $55,000 475 375 850 Totality of events 550 450 1,000 Male Total Income over $55,000 475 850 Totality of events 550 1,000

OpenStudy (amistre64):

The probability of picking a male, changes depending on the event we are describing. why??

OpenStudy (anonymous):

because of being male or female?

OpenStudy (amistre64):

no, it is becasue the tabled data gives us different probabilites for pick a man based on different events. the P(man) depends on the even we are confided to. P(man| over 55) = 475/850 P(man| all events) = 550/1000 since the probability of picking a man changes with the events, then it depends on which event we are concerned with.

OpenStudy (anonymous):

so P(being male | the person earns over $55,000) ≠ P(being male) .. correct?

OpenStudy (amistre64):

they are not equal, so the events are dependant lets do something simpler same setup but with independant stuff. m f 55+ 3 7 10 55- 27 63 90 30 70 100 P(m|55+) = 3/10 P(m|55-) = 27/90 P(m|all events) = 30/100 all three cases give us the same exact probability ... therefore the P(m) does not depend on any given event

OpenStudy (anonymous):

so A is correct?

OpenStudy (amistre64):

lol, yeah, if thats what we are taking from this A is fine

OpenStudy (anonymous):

yea just wanted to make sure .. thanks a lot for your patient